OpenStudy (anonymous):

d/dx {((X^2)+1)((X^2)+2)/((x^3)+1)((x^3)+2)} Is there an easier way to do this other than multiplying it all out and then using quotient rule?

OpenStudy (anonymous):

take logarithm

OpenStudy (anonymous):

Thats what I was thinking but I wasnt at class for that. How would taking the logarithm go exactly?

OpenStudy (anonymous):

\[\ln y=\ln \left( x^2+1 \right)+\ln \left( x^2+2 \right)-\ln \left( x^3+1 \right)-\ln \left( x^3+2 \right)\]

OpenStudy (anonymous):

\[\frac{ 1 }{ y }\frac{ dy }{ dx }=\frac{ 2x }{ x^2+1 }+?\]

OpenStudy (anonymous):

and then the derivative of each would simply be 1/whatever is in the parantheses

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

then take y to R.H.S and put the value of y

OpenStudy (anonymous):

Thank you